​A(n) 8.0%, ​25-year bond has a par value of​ $1,000 and a call price of $1,075.

​(The bond's first call date is in 5​ years.) Coupon payments are made semiannually​ (so use semiannual compounding where​ appropriate).

a. Find the current​ yield, YTM, and YTC on this​ issue, given that it is currently being priced in the market at $1,200.

Which of these 3 yields is the​ highest? Which is the​ lowest? Which yield would you use to value this​ bond? Explain.

b. Repeat the 3 calculations​ above, given that the bond is being priced at ​$850.

Now which yield is the​ highest? Which is the​ lowest? Which yield would you use to value this​ bond? Explain.
Answer #1

Solution :-

(a)

(i)

Price of Bond = $1,200

Annual Coupon = $1,000 * 8.0% = $80

Current Yield = $80 / $1,200 = 6.67%

(ii)

Time to Call = 5 Years

As Semiannual (n) = 5 * 2 = 10

Semiannual Coupon = $80 / 2 = $40

Fax ERATE(C3,C4,C5,C6) B с 1 C8 A (11 2 3 Time to Call 4 Coupon payment 5 Price 6 Price to Call 7 8 Yield to Call 9 10 11 Npe

Now Annual Yield to Call = 2.40% * 2 = 4.80%

(iii)

Time to Maturity = 25 Years

As Semiannual (n) = 25 * 2 = 50

12 (1) 50 Nper Pmt PV $40 ($1,200.00) $1,000 FV 13 14 Time to Maturity 15 Coupon payment 16 Price 17 Face Value 18 19 Yield t

Now Annual Yield to Maturity = 3.19% * 2 = 6.38%

Now Highest Yield is Current Yield

Lowest Yield is Yield to Call

To Value the Bond we use Yield to Maturity

(b)

(i)

Price of Bond = $850

Annual Coupon = $1,000 * 8.0% = $80

Current Yield = $80 / $850 = 9.41%

(ii)

Time to Call = 5 Years

As Semiannual (n) = 5 * 2 = 10

Semiannual Coupon = $80 / 2 = $40

fx =RATE(C3,C4,C5,C6) B с C8 A 1 ) 2 3 Time to Call 4 Coupon payment 5 Price 6 Price to Call Nper Pmt PV FV 10 $40 ($850.00)

Now Annual Yield to Call = 6.65% * 2 = 13.30%

(iii)

11 12 (1) 13 14 Time to Maturity 15 Coupon payment 16 Price 17 Face Value 18 19 Yield to Maturity 20 21 Nper Pmt PV FV 50 $40

Now Annual Yield to Maturity = 4.80% * 2 = 9.60%

Now Highest Yield is Yield to Call

Lowest Yield is Yield to Maturity

To Value the Bond we use Yield to Maturity

If there is any doubt please ask in comments

Thank you please rate